Calculate the forward discount of the euro against the dollar​ (the dollar is the home​ currency) if the spot rate is

$ 1.6257 divided by pound$1.6257/£

and the

66​-month

forward rate is

$ 1.6045 divided by pound$1.6045/£.

​Note: Use a​ 360-day year.

The forward premium on the dollar is

nothing​%
Answer #1

Spot Eur/Dollar = 1/1.6257 = Eur 0.6151/$

Forward Eur/Dollar = 1/1.6045 = Eur 0.6232/$

Forward premium = [(0.6232-0.6151)/0.6151]*(12/66)*100 = 0.239%